OpenStudy (anonymous):

in a right triangle, b=5, B=20, C=90, and A=70. solve for a and c. I tried using the Law of Sines, but I am very nearly sure I messed up somewhere because I keep getting no solution

OpenStudy (kc_kennylau):

sinB/b = sinC/c c = 5/sin(20) = 14.619

OpenStudy (kc_kennylau):

sinB/b = sinA/a a = 5sin(70)/sin(20) = 13.737

OpenStudy (anonymous):

ah, i see where i messed up now, thank you for clarifying. :)

OpenStudy (kc_kennylau):

glad to know you've solved your problem :)